OpenStudy (anonymous):

The area of a right triangle is 4. The sum of the lengths of the two sides adjacent to the right angle of the triangle is 9. What is the length of the hypotenuse of the triangle.

OpenStudy (asnaseer):



OpenStudy (asnaseer):

Just form some equations using the information that you are given.

OpenStudy (asnaseer):

e.g. area = \(\frac{1}{2}ab=4\)

OpenStudy (asnaseer):

and you are also told a+b=9

OpenStudy (asnaseer):

and you know from Pythagorus that \(c^2=a^2+b^2\)

OpenStudy (asnaseer):

do you need more help or do you know how to proceed from here?

OpenStudy (asnaseer):

do you understand?

OpenStudy (asnaseer):

Please let me know if you need any further help on this.

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

65 is apparently wrong

OpenStudy (asnaseer):

where did you get 65 from?

OpenStudy (anonymous):

that was my final answer

OpenStudy (asnaseer):

I think what you found was \(c^2=65\) - you now need to take its square root.

OpenStudy (anonymous):

you cant

OpenStudy (anonymous):

so u get +-65^(2)

OpenStudy (asnaseer):

the hypotenuse cannot be a negative number. so the answer should be:\[c^2=65\]therefore:\[c=\sqrt{65}\approx8.06\]

OpenStudy (anonymous):

ok thats the answer

OpenStudy (anonymous):

thanks

OpenStudy (asnaseer):

did you understand the process? that is more important than just getting the right answer.

OpenStudy (anonymous):

somewhat, yeah

OpenStudy (asnaseer):

ok - I'll be more than glad to explain it to you further if you wish.

OpenStudy (anonymous):

nah, im ok. thanks though. i appreciate it.

OpenStudy (asnaseer):

you are more than welcome - anytime :-)

OpenStudy (anonymous):

im gonna post some other questions i need help with though

OpenStudy (asnaseer):

ok